OpenStudy (anonymous):

What is 9/20 as a decimal? Please Help I Will Reward! ☺ ☻ ♥ ☺ ☻ ♥

OpenStudy (anonymous):

0.45 simple division

OpenStudy (anonymous):

Thx!

OpenStudy (luigi0210):

Or if you like multiplication.. \[\LARGE \frac{9}{20}*\frac{5}{5}=\frac{45}{100}=\frac{.45}{1}=0.45\]

OpenStudy (anonymous):

^ a littlebit more explained answer